OpenStudy (anonymous):

Amanda composed two songs. The longer song is 126 seconds less than twice the length of the short song. The longer song lasts 286 seconds. How long is the shorter song?

OpenStudy (anonymous):

L = 2s - 126 = 286 sec => s = (286 + 126 )/ 2 = 206 sec

OpenStudy (callisto):

Let x be the length of the shorter song From ''The longer song is 126 seconds less than twice the length of the short song'' length of longer song = 2x-126 = 286 So solve it from here, can you?

OpenStudy (callisto):

Oops... I'm too slow :S

OpenStudy (anonymous):

Sorry, I haven't noticed you're typing :/

OpenStudy (callisto):

^No sorry, not your problem though :)